INTRACAVITY LASER ABSORPTION SPECTROSCOPY OF ZIRCONIUM FLUORIDE IN THE NEAR INFRARED

\underline{LEAH C. O’BRIEN}, JACK C. HARMS, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652; JAMES J. O’BRIEN , Department of Chemistry and Biochemistry, University of Missouri, St Louis, MO 63121-4499.

A new band of ZrF has been recorded in the near-infrared with rotational resolution using intracavity laser absorption spectroscopy. A red-degraded bandhead is observed at 12527cm^{−1}, and 2 R-branches and 2 P-branches have been identified. The results of the analysis will be presented. The gas phase ZrF molecules were produced using a zirconiumlined hollow cathode in an argon-based electric discharge with a small amount of SF_{6}.